QUESTION 3 17) Let Xi. X. X be a random sample from a distribution with probability density function f(x, ?) | ße_ß, for x >0 elsewhere (a) What is the likelihood (LU) = L (x1.X2. xalß)) of the sample? Simplify it. (b) Use the factorization criterion/theorem to show that ? x, is a sufficient statistic for . 4
Answer #1

According to the random sample of X_{1},X_{2},.....X_{n} froma distribution with probability function ias

f(x)=\beta e^{-\beta x} for x\geq 0

=0 elsewhere .

(a) Now the likelihood estimation is calculated as

L(\beta\mid x )=\prod_{1}^{n}f(x,\beta )=\prod_{1}^{n}\beta e^{-\beta x}

  =\beta e^{-\beta x_{1}}\times \beta e^{-\beta x_{2}}\times \beta e^{-\beta x_{3}}\times ......\times \beta e^{-\beta x_{n}}

  =\beta^{n} e^{-\beta \sum_{1}^{n}x}

Taking log both side we get

logL(\beta\mid x )=log(\beta^{n} e^{-\beta \sum_{1}^{n}x})

  =nlog\beta -\beta \sum xloge......................(1)

Taking derivative both side with respect to \beta we get

\frac{dlogL(\beta\mid x )}{d\beta }=\frac{n}{\beta }-\sum x=0 first order condition of optimization.

or,\frac{n}{\beta }=\sum x

or,\beta =\frac{n}{\sum x}=\frac{1}{\bar{x}} as \bar{x}=\frac{\sum x}{n}

Therefore likelihood of the sample is \beta =\frac{1}{\bar{x}} (maximum likelihood estimator)

(b) According to the random sample of X_{1},X_{2},.....X_{n} froma distribution with probability function , joint probability density function is

f(x_{1},x_{2},.....x_{3},\beta )=f(x_{1},\beta )\times f(x_{2},\beta )\times f(x_{3},\beta )\times .....\times f(x_{n},\beta )

Therefore by replacing f(x)=\beta e^{-\beta x} we get

f(x_{1},x_{2},.....x_{3},\beta )=\beta e^{-\beta x_{1}}\times \beta e^{-\beta x_{2}}\times \beta e^{-\beta x_{3}}\times .....\times \beta e^{-\beta x_{n}}

f(x_{1},x_{2},.....x_{3},\beta )=\beta^{n} e^{-\beta\sum_{i=1}^{n} x_{i}

Now according to the factorization theorem we can factored the joint proabbility function as one which is depends on the parameter \beta with the statistics \sum_{i=1}^{n}X_{i} and another is independent on the parameter \beta as

f(x_{1},x_{2},.....x_{3},\beta )=[\beta^{n} e^{-\beta\sum_{i=1}^{n} x_{i}]\times 1

Where [\beta^{n} e^{-\beta\sum_{i=1}^{n} x_{i}] depends on parameter \beta and [1] is independent of parameter \beta .

Therefore the factorization theorem shows that  \sum_{i=1}^{n}X_{i}is a sufficient statistics for \beta and as \bar{x} is one to one relationship with \sum_{i=1}^{n}X_{i} , then \bar{x} also a sufficient statistics for \beta .